OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the third measure is the hypotenuse. 4,7, (radical sign)65

OpenStudy (hlambach):

Pythagorean Therorem: (leg1)^2 + (leg2)^2 = (hypotenuse)^2 Does 7^2 + 4^2 = sqrt(65)? If yes, then it can fit the sides of a right triangle.

OpenStudy (hlambach):

whoops: Does 7^2 + 4^2 = sqrt(65)^2

OpenStudy (abb0t):

pythatogean theorem = \(x^2 + y^2 = z^2\) where you have: